# CityEQA: A Hierarchical LLM Agent on Embodied Question Answering Benchmark in City Space

### v2

@@ Abstract
Embodied Question Answering (EQA) has primarily focused on indoor environments, leaving the complexities of urban settings—spanning environment, action, and perception—largely unexplored. To bridge this gap, we introduce CityEQA, a new task where an embodied agent answers open-vocabulary questions through active exploration in dynamic city spaces. To support this task, we present CityEQA-EC, the first benchmark dataset featuring 1,412 human-annotated tasks across six categories, grounded in a realistic 3D urban simulator. Moreover, we propose Planner-Manager-Actor (PMA), a novel agent tailored for CityEQA. PMA enables long-horizon planning and hierarchical task execution: the Planner breaks down the question answering into sub-tasks, the Manager maintains an object-centric cognitive map for spatial reasoning during the process control, and the specialized Actors handle navigation, exploration, and collection sub-tasks. Experiments demonstrate that PMA achieves 60.7% of human-level answering accuracy, significantly outperforming frontier-based baselines. While promising, the performance gap compared to humans highlights the need for enhanced visual reasoning in CityEQA. This work paves the way for future advancements in urban spatial intelligence. Dataset and code are available at https://github.com/tsinghua-fib-lab/CityEQA.git.
@@ 1 Introduction
Embodied Question Answering (EQA) Das et al. (2018) represents a challenging task at the intersection of natural language processing, computer vision, and robotics, where an embodied agent (e.g., a UAV) must actively explore its environment to answer questions posed in natural language. While most existing research has concentrated on indoor EQA tasks Gao et al. (2023); Peña-Narvaez et al. (2023), such as exploring and answering questions within confined spaces like homes or offices Liu et al. (2024a), relatively little attention has been dedicated to EQA tasks in open-ended city space. Nevertheless, extending EQA to city space is crucial for numerous real-world applications, including autonomous systems Kalinowska et al. (2023), urban region profiling Yan et al. (2024), and city planning Gao et al. (2024).
@@
1) Environmental complexity with ambiguous objects: Urban environments are inherently more complex, featuring a diverse range of objects and structures, many of which are visually similar and difficult to distinguish without detailed semantic information (e.g., buildings, roads, and vehicles). This complexity makes it challenging to construct task instructions and specify the desired information accurately, as shown in Figure 1.
@@
Place Open Vocab Active Platform Reference
EQA-v1 Indoor House3D Das et al. (2018)
IQUAD Indoor AI2-THOR Gordon et al. (2018)
MP3D-EQA Indoor Matterport3D Wijmans et al. (2019)
MT-EQA Indoor House3D Yu et al. (2019)
ScanQA Indoor - Azuma et al. (2022)
SQA3D Indoor - Ma et al. (2023)
K-EQA Indoor AI2-THOR Tan et al. (2023)
OpenEQA Indoor ScanNet/HM3D Majumdar et al. (2024)
CityEQA-EC City (Outdoor) EmbodiedCity -
@@
To address CityEQA tasks, we introduce the Planner-Manager-Actor (PMA), a novel baseline agent powered by large models, designed to emulate human-like rationale for solving long-horizon tasks in urban environments, as illustrated in Figure 1. PMA employs a hierarchical framework to generate actions and derive answers. The Planner module parses tasks and creates plans consisting of three sub-task types: navigation, exploration, and collection. The Manager oversees the execution of these plans while maintaining a global object-centric cognitive map Deng et al. (2024). This 2D grid-based representation enables precise object identification (retrieval) and efficient management of long-term landmark information. The Actor generates specific actions based on the Manager’s instructions through its components: Navigator, Explorer, and Collector. Notably, the Collector integrates a Multi-Modal Large Language Model (MM-LLM) as its Vision-Language-Action (VLA) module to refine observations and generate high-quality answers. PMA’s performance is assessed against four baselines, including humans. Results show that humans perform best in CityEQA, while PMA achieves 60.73% of human accuracy in answering questions, highlighting both the challenge and validity of the proposed benchmarks.
@@
  • Experimental results demonstrate that our approach outperforms existing baselines in tackling the CityEQA task. However, the gap with human performance highlights opportunities for future research to improve visual thinking and reasoning in embodied intelligence for city spaces.
Figure 2: Example questions and dataset statistics of CityEQA-EC.
@@ 2.1 Task Formulation
[…]
@@ 2.2 Dataset Collection and Validation
To obtain a high-quality dataset, we employed the EmbodiedCity Gao et al. (2024), which is a highly realistic 3D simulation platform based on the buildings, roads, and other elements in a real city. It is implemented using Unreal Engine 4 Sanders (2016) and Microsoft AirSim plugins Shah et al. (2018). The collection process is to determine the 4-tuple elements $\xi=(e,{{p}_{0}},q,y)$ of each instance. Unlike indoor simulators with many different scenes, EmbodiedCity is a coherent and extensive scene. As a result, for all instances, their scene $e$ corresponds to EmbodiedCity.
@@ 3.1 Overview
An overview of the proposed PMA agent for CityEQA tasks is shown in Figure 3. The PMA agent comprises three major modules: Planner, Manager, and Actor, all powered by pre-trained large models. The Planner is responsible for parsing the question $q$ and formulating an executable plan before any actions are taken. The Manager serves as the core module, receiving structured information from the Planner and processing observations at each time step to maintain an object-centric cognitive map using an MM-LLM. Additionally, through a process control module, the Manager issues task instructions to the Actor, which then utilizes various action generators to execute the required responses. Once the plan is completed, the Manager generates an answer based on its accumulated memory.
@@ 3.4 Actor Module
The typical exploration sub-task is described as Exploration(building_1, west, red_car), which means the goal is to explore the west side of building_1 to find a red car. The explorer uses the Move and Look Around (MLA) strategy due to the complexity of outdoor environments, where re-observing previously explored areas from different angles can yield different results. The exploration area is defined on the map based on landmark distribution and spatial relationships. A set of exploration points is generated within this area, maintaining a fixed distance of ${L^{exp}=10}$ meters between them. At each point, the agent thoroughly observes its surroundings by looking in four directions: front, back, left, and right. After completing observations at one point, the agent moves to the next closest point and continues until either the target object is found or all points are covered. A MM-LLM is employed to determine whether the target appears in any given observation.
@@
The collection sub-task instructions only include an information requirement. We provide an MM-LLM-driven Collector to gather the required information from observations. Additionally, the Collector can select an action from a predefined action set to fine-tune its observation view, enabling the collection of higher-quality information. More details of the design of Collector is presented in Appendix A.2.
@@ 4.1 Experiment setup
In CityEQA, we adopted three widely used metrics for evaluating EQA tasks Das et al. (2018): Question Answering Accuracy (QAA), Navigation Accuracy (NA), and Mean Time Step (MTS). QAA assesses the correctness of the answers by comparing them to the ground truth. The open-vocabulary nature of the CityEQA task poses challenges for evaluation. Inspired by OpenEQA Majumdar et al. (2024), we employed an LLM as the judge to assign scores $\theta\in\left\{{1,2,...,5}\right\}$ to the answers. For detailed information, please refer to the Appendix A.3. NA is measured by the distance between the agent’s final position and the target object ${p^{tar}}$ upon task completion, reflecting whether the agent successfully located and approached the target. MTS is calculated as the average number of time steps required to complete all tasks, indicating the efficiency of the embodied agent’s action strategy.
@@
We employed GPT-4o as the MM-LLM for visual analysis, which was utilized in both the Explorer and Collector modules. Meanwhile, GPT-4 was adopted as the text analysis model, responsible for question parsing, plan generation, answer generation, and automated scoring. For each task, the object-centric cognitive map is constructed centered around the agent’s initial pose, with a side length of 400 meters and a resolution of 1 meter. We considered buildings as landmarks and accounted for four spatial relationships: north, south, east, and west. Additionally, we limited the total number of time steps for navigation and exploration to 50 steps and restricted the number of steps for collection to 10 steps. Due to API limitations, 200 tasks are randomly selected from CityEQA-EC for the experiments.
@@
Our guiding principle is to investigate how to use foundation models to complete CityEQA tasks without any additional fine-tuning. Therefore, we employed four baselines that are widely employed in the studies of EQA tasks. More details of baselines can be found in Appendix A.3.
  • Blind Agents generate answers based solely on the text of questions without obtaining any visual inputs. It serves as a reference for assessing the extent to which one can rely purely on prior world knowledge and/or random guessing Majumdar et al. (2024).
  • LLM-VQA bypasse the active exploration process and is directly provided with the RGB image obtained from the ${p^{obs}}$ to answer the questions. This approach aims to assess the visual perception and reasoning capabilities of MM-LLMs in urban environments, while eliminating the interference of embodied actions.
  • Frontier-Based Exploration (FBE) Agent, commonly used indoor baseline Ren et al. (2024), does not utilize landmarks or spatial relationships.
  • Human Agents are employed to establish human-level performance metrics on our benchmark. We categorized human agents into two types, H-VQA and H-EQA. H-VQA is directly provided with an RGB image to perform Visual VQA tasks, similar to the setup of LLM-VQA. H-EQA launches from the initial pose and actively explores the environment based on the question description to find the answer.
4.2 Comparison with Baselines
Table 2: Performance of baselines and the proposed PMA on the CityEQA tasks. (PMA-7 means the PMA uses 7 steps to perform collection sub-tasks)
QAA (1-5) NA (m) MST
@@ 4.2 Comparison with Baselines
1. GPT-4 1.90±1.64 - -
2. Qwen-2.5 2.34±1.88 - -
LLM-VQA
1. GPT-4o 4.37±1.35 - -
2. Qwen-2.5 4.00±1.67 - -
FBE Agent 2.31±2.54 86.92±53.71 39.31±32.17
@@
1. H-VQA 4.87±0.72 - -
2. H-EQA 4.94±0.21 38.72±40.87 9.31±6.32
PMA-7 3.00±1.96 46.56±36.39 24.44±14.39
Figure 4: Categroy-level performance of the proposed PMA.
The results are shown in Table 2 and the category-level performance of PMA is shown in Figure 4. Some observations can be obtained:
  • The proposed PMA outperforms the Blind Agent and FBE Agent, as it leverages visual inputs and conducts more efficient perception activities guided by landmarks and spatial relationships. Compared to human agents, PMA shows a significant gap in QAA, achieving only 60.73% of H-EQA. However, despite the considerable difference in MST, the NA gap is relatively small. This reveals that PMA’s navigation and exploration strategies are effective, allowing it to approach target objects even with more time steps.
  • PMA’s performance varies across task types. It achieves the highest QAA on World Knowledge tasks, likely because these tasks rely partially on the LLM’s inherent knowledge and require minimal visual inputs. However, it performs the worst on Object Recognition tasks due to their open-ended answers and greater reliance on visual inputs.
  • Humans excel in both H-VQA and H-EQA tasks. Notably, the QAA of H-EQA is slightly higher than that of H-VQA, indicating actively adjusting the observation view helps address challenges like occlusion and reflection. An illustrative case is provided in Appendix A.3.
  • The FBE Agent performs poorly, with a QAA even lower than that of the blind Qwen2.5. This highlights the importance of utilizing landmarks and spatial relationships in exploring urban environments. It also indicates that embodied models designed for indoor environments cannot be directly applied to open-ended city space.
  • LLM-VQA correctly answers most questions, although its QAA is lower than humans. This confirms the validity of our dataset. Moreover, the performance gap between Qwen-2.5 and GPT-4o indicates that the inherent differences in visual understanding and reasoning capabilities of MM-LLMs are also important factors influencing agent performance.
  • The Blind Agent achieves a certain level of accuracy, although it is significantly lower than that of humans and GPT-4o. This reveals the regularities of the real world that can be leveraged for answering questions.
@@
4.3 Study on Collector Module
Previous experimental results have confirmed the effectiveness of navigation and exploration strategies in PMA. Therefore, in this section, we aim to investigate the impact of fine-grained adjustments in observations on performance. To achieve this, we recorded the Collector’s pose at each step (up to 10 steps) along with the generated responses and calculated relevant metrics. The results are shown in Figure 5.

### v3

@@ Abstract
Embodied Question Answering (EQA) has primarily focused on indoor environments, leaving the complexities of urban settings—spanning environment, action, and perception—largely unexplored. To bridge this gap, we introduce CityEQA, a new task where an embodied agent answers open-vocabulary questions through active exploration in dynamic city spaces. To support this task, we present CityEQA-EC, the first benchmark dataset featuring 1,412 human-annotated tasks across six categories, grounded in a realistic 3D urban simulator. Moreover, we propose Planner-Manager-Actor (PMA), a novel agent tailored for CityEQA. PMA enables long-horizon planning and hierarchical task execution: the Planner breaks down the question answering into sub-tasks, the Manager maintains an object-centric cognitive map for spatial reasoning during the process control, and the specialized Actors handle navigation, exploration, and collection sub-tasks. Experiments demonstrate that PMA achieves 60.7% of human-level answering accuracy, significantly outperforming competitive baselines. While promising, the performance gap compared to humans highlights the need for enhanced visual reasoning in CityEQA. This work paves the way for future advancements in urban spatial intelligence. Dataset and code are available at https://anonymous.4open.science/r/CityEQA-3027.
@@ 1 Introduction
Embodied Question Answering (EQA) Das et al. (2018) represents a challenging task at the intersection of natural language processing, computer vision, and robotics, where an embodied agent (e.g., a UAV) must actively explore its environment to answer questions posed in natural language. While most existing research has concentrated on indoor EQA tasks Gao et al. (2023); Peña-Narvaez et al. (2023) or traditional indoor/outdoor Visual Question Answering (VQA) tasks Sun et al. (2024), relatively little attention has been dedicated to EQA tasks in open-ended city space, as shown in Table 1. Nevertheless, extending EQA to city space is crucial for numerous real-world applications, including autonomous systems Kalinowska et al. (2023), urban region profiling Yan et al. (2024), and city planning Gao et al. (2024).
@@
1) Environmental complexity with ambiguous objects: Urban environments are inherently more complex, featuring a diverse range of objects and structures, many of which are visually similar and difficult to distinguish without detailed semantic information (e.g., buildings, roads, and vehicles). This complexity makes it challenging to construct task instructions and specify the desired information accurately Ji et al. (2025); Xu et al. (2025).
@@
Platform Reference Place Open Vocab Active
EQA-v1 House3D Das et al. (2018) Indoor
IQUAD AI2-THOR Gordon et al. (2018) Indoor
MP3D-EQA Matterport3D Wijmans et al. (2019) Indoor
MT-EQA House3D Yu et al. (2019) Indoor
K-EQA AI2-THOR Tan et al. (2023) Indoor
HM-EQA HM3D Ren et al. (2024) Indoor
S-EQA VirtualHome Dorbala et al. (2024) Indoor
NoisyEQA - Wu et al. (2024) Indoor
OpenEQA ScanNet/HM3D Majumdar et al. (2024) Indoor
City-3DQA - Sun et al. (2024) Outdoor
EarthVQA - Wang et al. (2024) Outdoor
Open3DVQA - Zhan et al. (2025) Outdoor
CityEQA-EC EmbodiedCity - Outdoor
@@
To address CityEQA tasks, we introduce the Planner-Manager-Actor (PMA), a novel baseline agent powered by large models, designed to emulate human-like rationale for solving long-horizon tasks in urban environments, as illustrated in Figure 1. PMA employs a hierarchical framework to generate actions and derive answers. The Planner module parses tasks and creates plans consisting of three sub-task types: navigation, exploration, and collection. The Manager oversees the execution of these plans while maintaining a global object-centric cognitive map Deng et al. (2024). This 2D grid-based representation enables precise object identification (retrieval) and efficient management of long-term landmark information. The Actor generates specific actions based on the Manager’s instructions through its components: Navigator, Explorer, and Collector. Notably, the Collector integrates the Vision Language Model (VLM) as its Vision Language Action (VLA) module to refine observations and generate high-quality answers. PMA’s performance is assessed against five types of baselines, including humans. Results show that humans perform best in CityEQA, while PMA achieves 60.73% of human accuracy in answering questions, highlighting both the challenge and validity of the proposed benchmarks.
Figure 2: Task examples and dataset statistics of the CityEQA-EC.
@@
  • Experimental results demonstrate that our approach outperforms existing baselines in tackling the CityEQA task. However, the gap with human performance highlights opportunities for future research to improve visual thinking and reasoning in embodied agents for city spaces.
@@ 2.1 Task Formulation
An instance of the EQA task is defined by the 4-tuple: $\xi=(e,q,y,{{p}_{0}})$, where $e$ is the simulated or real 3D scene that agent can interact with, $q$ is the question, and $y$ is the ground truth answer. The ${{p}_{0}}$ denotes the agent’s initial pose, including 3D position and orientation. Given the instance $\xi$, the goal is for the embodied agent (e.g., drones) to complete the task by gathering the required information from $e$ and generating the answer $\hat{y}$ in response to $q$. Specifically, the agent starts at the initial pose ${{p}_{0}}$ and interacts with the scene $e$ step by step. At each time step $t$, the agent can move to a specific pose ${{p}_{t}}$, and obtain an observation ${{o}_{t}}=(I_{t}^{rgb},I_{t}^{d})$ from the scene, where $I_{t}^{rgb}\in{{\mathbb{R}}^{H\times W\times 3}}$ is the RGB image and $I_{t}^{d}\in{{\mathbb{R}}^{H\times W}}$ is the depth image. Based on these observations, the agent generates the answer $\hat{y}$. The key challenge is to produce a high-quality answer while minimizing the time steps required.
@@ 2.2 Dataset Collection and Validation
To obtain a high-quality dataset, we employed EmbodiedCity Gao et al. (2024), which is a highly realistic 3D simulation platform based on the buildings, roads, and other elements in a real city. It is implemented using Unreal Engine 4 Sanders (2016) and Microsoft AirSim plugins Shah et al. (2018). The collection process is to determine the 4-tuple elements $\xi=(e,{{p}_{0}},q,y)$ of each instance. Unlike indoor simulators with many different scenes, EmbodiedCity is a coherent and extensive scene. As a result, for all instances, their scene $e$ corresponds to EmbodiedCity.
@@ 3.1 Overview
An overview of the proposed PMA agent for CityEQA tasks is shown in Figure 3. The PMA comprises three major modules: Planner, Manager, and Actor, all powered by pre-trained foundation models. Planner is responsible for parsing the question $q$ and formulating an executable plan before any actions are taken. Manager serves as the core module, receiving structured information from Planner and processing observations at each time step to maintain an object-centric cognitive map using an VLM. Additionally, through a process control module, Manager issues task instructions to Actor, which then utilizes various action generators to execute the required responses. Once the plan is completed, Manager generates an answer based on its accumulated memory.
@@ 3.4 Actor Module
The typical exploration sub-task is described as Exploration(building_1, west, red_car), which means the goal is to explore the west side of building_1 to find a red car. The explorer uses the Move and Look Around strategy due to the complexity of outdoor environments, where re-observing previously explored areas from different angles can yield different results. The exploration area is defined on the map based on landmark distribution and spatial relationships. A set of exploration points is generated within this area, maintaining a fixed distance of ${L^{exp}=10}$ meters between them. At each point, the agent thoroughly observes its surroundings by looking in four directions: front, back, left, and right. After completing observations at one point, the agent moves to the next closest point and continues until either the target object is found or all points are covered. A VLM is employed to determine whether the target appears in any given observation.
@@
The collection sub-task instructions only include an information requirement. We provide a VLM-driven Collector to gather the required information from observations. Additionally, the Collector can select an action from a predefined action set to fine-tune its observation view, enabling the collection of higher-quality information. More details of Collector is presented in Appendix A.2.
@@ 4.1 Experiment setup
In CityEQA, we adopted three widely used metrics for evaluating EQA tasks Das et al. (2018): Question Answering Accuracy (QAA) assesses the correctness of the answers by comparing them to the ground truth. The open-vocabulary nature of the CityEQA task poses challenges for evaluation. Inspired by OpenEQA Majumdar et al. (2024), we employed an LLM as the judge to assign scores $\theta\in\left\{{1,2,...,5}\right\}$ to the answers. For detailed information, please refer to the Appendix A.3. Navigation Error (NE) is measured by the distance between the agent’s final position and the target object ${p^{tar}}$ upon task completion, reflecting whether the agent successfully located and approached the target. Mean Time Step (MTS) is calculated as the average number of time steps required to complete all tasks, indicating the efficiency of the embodied agent’s action strategy.
@@
For each task, the object-centric cognitive map is constructed centered around the agent’s initial pose, with a side length of 400 meters and a resolution of 1 meter. The dimension of the images obtained by the agent is 640×480, and we considered buildings as landmarks and accounted for four spatial relationships: north, south, east, and west. Additionally, the total number of time steps for navigation and exploration is limited to 50 steps and the maximum steps for collection is 10. GPT-4o and GPT-4 are the default VLM and LLM used in the PMA. Due to API limitations, 200 tasks are randomly selected from CityEQA-EC for the experiments.
@@
We compare various models in a zero-shot setting, including five categories of baselines that are widely used in studies of EQA tasks. More details of baselines can be found in Appendix A.3.
  • Blind Agents Majumdar et al. (2024) generate answers based solely on the text of questions without obtaining any visual inputs. It serves as a reference for assessing the extent to which one can rely purely on prior world knowledge and/or random guessing.
  • Socratic Agents Jiang et al. (2025) use the VLM (GPT-4o) to convert the visual input during the exploration process into image captions, and then uses LLMs to generate answers based on these descriptions.
  • VQA Agents bypass the active exploration process and is directly provided with the RGB image obtained from the ${p^{obs}}$ to answer the questions. This approach aims to assess the visual perception and reasoning capabilities of VLMs in urban environments, while eliminating the interference of embodied actions.
  • Exploring Agents Ren et al. (2024) actively acquire visual inputs using Random Exploration (RE) and Frontier-Based Exploration (FBE), both commonly used as indoor baselines.
  • Human Agents are employed to establish human-level performance metrics on our benchmark. We categorize human agents as H-VQA or H-EQA, depending on whether they actively acquire visual inputs.
4.2 Comparison with State-of-the-art
Table 2: Performance of baselines and the proposed PMA on the CityEQA tasks.
QAA (1-5) $\uparrow$ NE (m) $\downarrow$ MTS $\downarrow$
@@ 4.2 Comparison with State-of-the-art
GPT-4 1.90±1.64 - -
Qwen-2.5 2.34±1.88 - -
LLaMA-v3.1-8b 2.31±1.72 - -
DeepSeek-v3 2.03±1.41 - -
Socratic Agents (VLM/LLM)
GPT-4o/GPT-4 2.71±1.72 - -
GPT-4o/Qwen-2.5 2.77±1.49 - -
GPT-4o/LLaMA-v3.1-8b 2.70±1.71 - -
GPT-4o/DeepSeek-v3 2.82±1.53 - -
VQA Agents
GPT-4o 4.37±1.35 - -
Qwen-2.5 4.00±1.67 - -
LLaVA-v1.5-7b 3.81±2.01 - -
Exploring Agents
RE 2.19±2.64 73.31±45.43 46.41±10.41
FEB 2.31±2.54 86.92±53.71 39.31±32.17
@@
H-VQA 4.87±0.72 - -
H-EQA 4.94±0.21 38.72±40.87 9.31±6.32
PMA (ours) 3.00±1.96 46.56±36.39 24.44±14.39
As shown in Table 2, human agents in both VQA and EQA settings achieve the highest QAA scores—4.87±0.72 for H-VQA and 4.94±0.21 for H-EQA—representing the upper bound for answer quality. They also demonstrate exceptional efficiency, with the lowest navigation error (38.72±40.17m) and completion steps (9.31±6.32), setting the gold standard for both quality and efficiency.
For automated methods, VQA agents like GPT-4o reach QAA scores up to 4.37±1.35, approaching human performance in answer quality, but lack active exploration abilities, preventing assessment of their overall task efficiency. Blind and Socratic agents perform significantly worse, with QAA between 1.90 and 2.82, showing the shortcomings of methods without visual information or with only language-based reasoning.
Exploring agents such as RE and FEB can handle active exploration and answering, but their QAA scores are low (2.19–2.31) and their NA and MTS are much higher (e.g., FEB: 86.92±53.71m, 39.31±32.17), resulting in less effective execution. In contrast, PMA achieves a balanced performance: its QAA of 3.00±1.96 is higher than all exploring, blind, and Socratic agents, though still just 60.73% that of H-EQA. Importantly, PMA’s navigation error (46.56±36.39m) and completion step (24.44±14.39) are dramatically less than traditional exploring agents, demonstrating notable practical gains.
@@
4.3 Ablation Studies
We conduct ablation studies on the Object-Centric Cognitive Map, navigator, and explorer modules in PMA, as shown in Table 3. Removing any of these modules leads to a significant decline in performance. Without the map, the agent becomes confused by similar landmarks in the environment and fails to perform effective active perception, resulting in the worst ablation outcome. Furthermore, the absence of the navigator is more detrimental than that of the explorer, further highlighting the importance of landmark-based navigation in urban environments.
Table 3: Ablation results.
QAA (1-5) $\uparrow$ NE (m) $\downarrow$ MTS $\downarrow$
PMA w/o map 2.31±1.82 76.41±48.64 43.27±31.92
PMA w/o navigator 2.33±1.64 68.31±46.91 38.83±27.71
PMA w/o explorer 2.68±1.87 57.13±41.43 20.62±15.11
PMA 3.00±1.96 46.56±36.39 24.44±14.39
4.4 Effectiveness of Collector Module
This section further investigates the effectiveness of the collector module, specifically the impact of fine-grained observation adjustments on performance. We recorded the observation at each step (10 steps in total) during the collection phase and calculated relevant metrics, as shown in Figure 4.